OpenStudy (anonymous):

how can i solve and find the polynomial f(x) that has the roots of –2, 5 of multiplicity 2. Explain how you would verify the zeros of f(x)

OpenStudy (amistre64):

write it up in factored form

OpenStudy (amistre64):

-2 5 5 (-2+2)(5-5)(5-5) = 0 (x+2)(x-5)(x-5) = 0

OpenStudy (anonymous):

\[(x-2)(x+5)^{2}\]

OpenStudy (amistre64):

close, they give you the xs; and say undo it

OpenStudy (amistre64):

-2-2 not= 0

OpenStudy (amistre64):

5+5 not= 0

OpenStudy (anonymous):

so it would be (x-2)(2x-10x+25) ?

OpenStudy (amistre64):

not quite

OpenStudy (amistre64):

x-2 is STILL wrong

OpenStudy (anonymous):

x+2

OpenStudy (amistre64):

and ill assume 2x meant x^2 :)

OpenStudy (anonymous):

yeah lol

OpenStudy (amistre64):

(x+2)(x^2-10x+25) good, now keep going :)

OpenStudy (anonymous):

\[x ^{3}-8x ^{2}+5x+50\] ?

OpenStudy (amistre64):

x^3-10x^2+25x 2x^2-20x+50 -------------------- yep, you got it

OpenStudy (anonymous):

so how would i verify the zeros?

OpenStudy (amistre64):

plug in the roots and see if you get zeros back :)

OpenStudy (amistre64):

synth it if you want

OpenStudy (anonymous):

how would i set that up?

OpenStudy (amistre64):

the confusing way apparently lol

OpenStudy (amistre64):

just like last question, pull the coeffs and synth it

OpenStudy (anonymous):